# Christian Ghiglino    Alastair Langtry

### 28 Jan 2024 v3

@@ Abstract
This paper adapts ideas from social identity theory to set out a new framework for modelling conspicuous consumption. Notably, this approach can explain two stylised facts about conspicuous consumption that initially seem at odds with one another, and to date have required different families of models to explain each: (1) people consume more visible goods when their neighbours’ incomes rise, but (2) consume less visible goods when incomes of those with the same race in a wider geographic area rise. The first fact is typically explained by ‘Keeping up with the Joneses’ models, and the second by signalling models. Our model also explains related features of conspicuous consumption: that the rich are more sensitive to others’ incomes than the poor, and that the effect of income inequality on consumption differs qualitatively across groups. Importantly, it explains this fourth stylised fact without falling back on differences in preferences across groups, as required in other models. In addition, our model delivers new testable predictions regarding the role of network structure and income inequality for conspicuous consumption.
@@
income inequality
Economists widely recognise that people’s consumption decisions respond to the consumption decisions of others. Existing literature documents two main stylised facts about these responses. First, an influential paper by Charles et al. (2009) documents that, holding their own permanent income fixed, people in the US consume less when average income of those of the same race and who live in the same state is higher. Second, work by Kuhn et al. (2011) shows that people consume more when average income of their geographic neighbours increases. This is often called a “Keeping up with The Joneses” effect. On the face of it, these two stylised facts may appear to conflict. Plus, the classes of models typically used to rationalise each of them – signalling models for the Charles et al. (2009) result and social comparisons models for the Kuhn et al. (2011) result – cannot capture both simultaneously.
We present a simple model of social identity that explains both of these stylised facts as a single phenomenon. In it, people derive status directly from consuming visible goods themselves, and from belonging to an identity that consumes lots of visible goods. In our model, identity group status depends on average consumption of visible goods by members of the identity group, and we assume there is some substitutability between status from each of these two sources. People also feel pressure to conform to some benchmark for their identity, and we assume it depends only on a person’s neighbours rather than their entire identity group. These two features are both core components of social identity theory (Costa-Font and Cowell, 2015). Our model is also able to speak to the first stylised fact without relying on the characteristics in the lowest part of the distribution – which is needed in signalling models – aligning our model and its predictions more closely with available data.
@@
Which of the two forces wins out depends on how closely connected the single person is to the people whose incomes increase. This is precisely what delivers both of the stylised facts. When the people whose incomes increase are close neighbours (as in Kuhn et al. (2011)) we should expect the second force to win out – so she consumes more visible goods. In contrast, when these people are not closely connected to her (as in Charles et al. (2009)) we should expect the first force to win out – delivering the seemingly counter-intuitive result that she consumes fewer visible goods.
@@ 1 Literature
Our paper sits within a large literature on interpersonal comparisons and conspicuous consumption. [^4] It spans disciplines, including marketing, sociology, social psychology (see Dubois and Ordabayeva (2017) and Dubois et al. (2021) for recent reviews), and economics (where it dates at least to Veblen (1899), with other early contributions from Duesenberry (1949); Festinger (1954); Frank (1985b, a) and Van de Stadt et al. (1985)). Bursztyn and Jensen (2017) provide an interesting survey of the wider topic of social pressure. Within economics, there are two broad views on why people engage in conspicuous consumption. The first assumes that people care about their income or wealth relative to others – and conspicuous consumption is used as a signal of high income/wealth. Charles et al. (2009) use this signalling approach to explain the first stylised fact – when others get richer, a given agent moves down the income ranking, and so has fewer people she needs to signal she is richer than (their model follows Glazer and Konrad (1996) very closely). However, with no explicit notion of ‘neighbours’, these models struggle to explain the second stylised fact – that people consume more when their neighbours’ incomes rise. Additionally, because agents are consuming to differentiate themselves from people poorer than themselves, everyone’s consumption decisions can be sensitive to the income of those at the very bottom of the distribution.
@@
The second view assumes that people make social comparisons direction over consumption – so people gain status from conspicuous consumption itself. The core idea is that people like high consumption relative to others, not just in absolute terms – often called a “Keeping up with the Joneses” motive. It is perhaps the more dominant of the two views. Our paper is most closely related to a recent strand of this literature, started by Ghiglino and Goyal (2010), that uses explicit networks to allow for very rich patterns of social comparisons. More recent contributions that use explicit networks include Immorlica et al. (2017); Langtry (2022) and Bramoullé and Ghiglino (2022). [^5] Before explicit networks were introduced, most authors either assumed that people care about their consumption relative to the population average (Abel, 1990; Clark and Oswald, 1996; Ljungqvist and Uhlig, 2000) or about their rank within the population (Frank, 1985a, b, 2001; Hopkins and Kornienko, 2004). More recent contributions include Luttmer (2005); Heffetz (2011); Frank et al. (2014); Drechsel-Grau and Schmid (2014); Alvarez-Cuadrado et al. (2016); Jinkins (2016); Bertrand and Morse (2016) and De Giorgi et al. (2020). These ‘Keeping up with the Joneses’ models successfully explain the second stylised fact. But they cannot explain the first. This is because they embed a desire to ‘Keep up with the Joneses’, which makes agents’ consumption choices strategic complements – an increase in some people’s income must (weakly) increase everyone’s consumption.
What our paper adds is the ability to explain a set of so-far difficult to reconcile stylised facts in a single framework. It does so by building a new type of model – one motivated by social identity, rather than signalling or ‘Keeping up with the Joneses’ motives.
@@
In doing so, we also relate to a strand of literature that examines the link between inequality and conspicuous consumption. By and large, formal models provide ambiguous results on the nature of the link. In Charles et al. (2009), a regressive redistribution of income (from a poorer agent to a richer one) increases total conspicuous consumption only if consumption is convex in income – but this is not a consistent property of observed consumption data (Heffetz, 2011). [^6] Other versions of the signalling model make similar predictions (Ireland, 2001; Heffetz, 2011; Moav and and Neeman, 2012; Jinkins, 2016). Further, Bilancini and Boncinelli (2012) show that the relationship between conspicuous consumption and inequality can be sensitive to the choice of whether to use ordinal or cardinal utility (also see Hopkins (2023) for similar discussion).
@@
In line with this literature, we do not present an ironclad relationship between inequality and conspicuous consumption. But we do provide novel predictions regarding how the nature of the relationship depends on properties of the network. Notably, it allows us to explain our fourth stylised fact – that higher dispersion in income within a person’s reference group leads to lower conspicuous consumption among Whites, but higher conspicuous consumption among Blacks – without an appeal to differences in underlying preferences across groups, but instead different network structures across groups.
@@
Our main contribution here is to show how models of social identity can help resolve a puzzle regarding conspicuous consumption behaviour – something these models have not been used for to date. From a more technical standpoint, we also make the benchmark that agents find it costly to deviate from (often called a ‘prototype’ in the literature) endogenous. In effect, this demands that there are not persistent differences between actual behaviour in equilibrium and the benchmark for ‘normal’ behaviour. Our utility formulation closely follows Shayo (2009) but with an endogenous ‘prototype’. However, we abstract away from the choice of identity group – which is often a key focus in this literature. This is both in the interest of tractability and because it is not needed to explain the stylised facts we focus on.
@@ 2 Model
with $\alpha>0$, $\beta>0$, $\gamma>0$. This strong functional form assumption allows for more transparent analysis, especially regarding the role of the network.
@@ 2.2 Discussion
Note that in line with typical social identity models, agents dislike being away from the reference point here – either above or below. This takes the view that people want to conform. It contrasts with the standard assumption in the social comparisons literature that people want to get ahead – and so dislike being below the reference point, but enjoy being above it. Nevertheless, it is possible to replace the social identity style desire to conform with a social comparisons style desire to get ahead. This change does not alter the key insights or the qualitative features of equilibrium behaviour. We show this in the Online Appendix. [^10] To remain highly tractable, it requires a linear, rather than quadratic, cost of consumption.
@@ 4 The effects of changes in income
    In the US, people consume less visible goods when the average income of others of the same race in their state is higher (Charles et al., 2009).
@@ 5.3 Explaining an empirical feature
We can rationalise these qualitatively different responses to changing inequality across racial groups as perhaps being due to a non-monotonic relationship between community density and community income. More precisely, suppose that density (a la 4) follows a U-shaped pattern: it is first decreasing and then increasing in income. Note that density relates to how people set the ‘benchmark for normal’ that they want to conform to, and hence how strongly people make social comparisons. This non-monotonicity essentially assumes that low and high income agents/communities feel the strongest pressures to conform.
Under this assumption, an increase in inequality among only high income communities will involve shifting income from lower density communities to higher density ones. 3(iii) tells us that consumption in the communities whose income did not change will then rise. Conversely, an increase in inequality among only low income communities will involve shifting income from higher density communities to lower density ones. In this case, 3(iii) tells us that consumption in the communities whose income did not change will then fall. Figure 1 provides an illustration.
@@
If such a non-monotonic relationship between density and income exists, then we should reasonably expect Blacks to be disproportionately below the peak of the curve, and Whites to be disproportionately above it. This is simply due to the well-documented fact that Blacks have, on average, lower incomes than White (Charles et al., 2009). So an increase in inequality among White would raise consumption of visible goods by Whites (among those whose income did not change), and the reverse would happen for Blacks. This matches exactly the finding by Charles et al. (2009).
Given that (often strong) homophily along many different dimensions is an extremely robust finding in the empirical literature on social networks (Currarini et al. (2009),McPherson et al. (2001),Currarini and Mengel (2016), Jackson et al. (2022)), we contend it is plausible that agents make stronger comparisons with those who are more similar to themselves in terms of location and income. If so, then the non-monotonic relationship that we posit in Figure 1 could be consistent with a pair of known stylised facts.
First, Bailey et al. (2020) show that higher income agents have more geographically dispersed social networks – proportionately less of their friends are nearby. Second, Chetty et al. (2022) show that there is substantial homophily based on socioeconomic status (which is closely tied to income) – people are disproportionately friends with those who have similar income to themselves. Importantly, they find that this homophily is substantially stronger at the upper end of the income distribution. Within the confines of our model, this would push richer agents to have stronger social comparisons, and so higher centrality.

### 28 Feb 2024 v4

@@ Abstract
This paper adapts ideas from social identity theory to set out a new framework for modelling conspicuous consumption. Agents derive status from their own conspicuous consumption and from belonging to an identity group with high conspicuous consumption. Importantly, these two sources of status are substitutes. Agents also feel pressure to conform with their neighbours in a network. This framework can rationalise a set of seemingly conflicting stylised facts about conspicuous consumption that are currently explained by different families of models. In addition, our model delivers new testable predictions regarding the effect of network structure and income inequality on conspicuous consumption.
@@
conspicuous consumption, income inequality
Economists widely recognise that people’s consumption decisions respond to the consumption decisions of others. Existing literature documents two main stylised facts about these responses. First, an influential paper by Charles et al. (2009) documents that, holding their own permanent income fixed, people in the US consume less when average income of those of the same race and who live in the same state is higher. Khamis et al. (2012) find similar behaviour for caste/religious groups in India. Second, work by Kuhn et al. (2011) shows that people consume more when average income of their geographic neighbours increases. This is often called a “Keeping up with The Joneses” effect. On the face of it, these two stylised facts may appear to conflict. Plus, the classes of models typically used to rationalise each of them – signalling models for the Charles et al. (2009) result and social comparisons models for the Kuhn et al. (2011) result – cannot capture both simultaneously.
We present a simple model of social identity that explains both of these stylised facts as a single phenomenon. In it, people derive status directly from consuming visible goods themselves, and from belonging to an identity that consumes lots of visible goods. They may also get intrinsic benefits from their own consumption, but adding this has no impact on the insights or the outcomes. In our model, identity group status depends on average consumption of visible goods by members of the identity group, and we assume there is some substitutability between status from each of these two sources. People also feel pressure to conform to some benchmark for their identity, and we assume it depends only on a person’s neighbours rather than their entire identity group. These two features are both core components of social identity theory (Costa-Font and Cowell, 2015). Our model is also able to speak to the first stylised fact without relying on the characteristics in the lowest part of the distribution – which is needed in signalling models – aligning our model and its predictions more closely with available data.
@@
Which of the two forces wins out depends on how closely connected the single person is to the people whose incomes increase. This is precisely what delivers both of the stylised facts. When the people whose incomes increase are close neighbours (as in Kuhn et al. (2011) and Khamis et al. (2012)) we should expect the second force to win out – so she consumes more visible goods. In contrast, when these people are not closely connected to her (as in Charles et al. (2009)) we should expect the first force to win out – delivering the seemingly counter-intuitive result that she consumes fewer visible goods.
@@ 1 Literature
Our paper sits within a large literature on interpersonal comparisons and conspicuous consumption. [^4] It spans disciplines, including marketing, sociology, social psychology (see Dubois and Ordabayeva (2017) and Dubois et al. (2021) for recent reviews), and economics (where it dates at least to Veblen (1899), with other early contributions from Duesenberry (1949); Festinger (1954); Frank (1985b, a) and Van de Stadt et al. (1985)). Bursztyn and Jensen (2017) provide an interesting survey of the wider topic of social pressure. Within economics, there are two broad views on why people engage in conspicuous consumption. The first assumes that people care about their income or wealth relative to others – and conspicuous consumption is used as a signal of high income/wealth. Charles et al. (2009) and Khamis et al. (2012) use this signalling approach to explain the first stylised fact – when others get richer, a given agent moves down the income ranking, and so has fewer people she needs to signal she is richer than (their model follows Glazer and Konrad (1996) very closely). However, with no explicit notion of ‘neighbours’, these models struggle to explain the second stylised fact – that people consume more when their neighbours’ incomes rise. Additionally, because agents are consuming to differentiate themselves from people poorer than themselves, everyone’s consumption decisions can be sensitive to the income of those at the very bottom of the distribution.
@@
The second view assumes that people make social comparisons direction over consumption – so people gain status from conspicuous consumption itself. The core idea is that people like high consumption relative to others, not just in absolute terms – often called a “Keeping up with the Joneses” motive. It is perhaps the more dominant of the two views. Earlier work assumed people either care about their consumption relative to the population average, or to their rank within the population. Important contributions to these two approaches include Abel (1990); Clark and Oswald (1996); Ljungqvist and Uhlig (2000), and (Frank, 1985a, b, 2001; Hopkins and Kornienko, 2004) respectively. [^5] More recent contributions include Luttmer (2005); Heffetz (2011); Frank et al. (2014); Drechsel-Grau and Schmid (2014); Alvarez-Cuadrado et al. (2016); Jinkins (2016); Bertrand and Morse (2016),De Giorgi et al. (2020) and Hopkins (2023) Our paper is most closely related to a recent strand within this literature, started by Ghiglino and Goyal (2010), that uses explicit networks to allow for very rich patterns of social comparisons (more recent contributions to this strand include Immorlica et al. (2017); Langtry (2022) and Bramoullé and Ghiglino (2022)). These ‘Keeping up with the Joneses’ models successfully explain the second stylised fact. But they cannot explain the first. This is because they embed a desire to ‘Keep up with the Joneses’, which makes agents’ consumption choices strategic complements – an increase in some people’s income must (weakly) increase everyone’s consumption.
@@
What our paper adds is the ability to explain a set of so-far difficult to reconcile stylised facts in a single framework. It does so by building a new type of model – one motivated by social identity, rather than signalling or ‘Keeping up with the Joneses’ motives. In doing so, we also relate to a strand of literature that examines the link between inequality and conspicuous consumption. By and large, formal models provide ambiguous results on the nature of the link. In Charles et al. (2009), a regressive redistribution of income (from a poorer agent to a richer one) increases total conspicuous consumption only if consumption is convex in income – but this is not a consistent property of observed consumption data (Heffetz, 2011). [^6] Other versions of the signalling model make similar predictions (Ireland, 2001; Heffetz, 2011; Moav and and Neeman, 2012; Jinkins, 2016). Further, Bilancini and Boncinelli (2012) show that the relationship between conspicuous consumption and inequality can be sensitive to the choice of whether to use ordinal or cardinal utility (also see Hopkins (2023) for similar discussion).
@@
In line with this literature, we do not present an ironclad relationship between inequality and conspicuous consumption. But we do provide novel predictions regarding how the nature of the relationship depends on properties of the network. Notably, it allows us to explain our fourth stylised fact – that higher dispersion in income within a person’s reference group leads to lower conspicuous consumption among Whites, but higher conspicuous consumption among Blacks. It does so without an appeal to differences in underlying preferences across groups, but instead relies on well-documented differences in average income.
@@
Our main contribution here is to show how models of social identity can help resolve a puzzle regarding conspicuous consumption behaviour – something these models have not been used for to date. From a more technical standpoint, we also make the benchmark that agents find it costly to deviate from (often called a ‘prototype’ in the literature) endogenous. Aside from this, our utility formulation closely follows Shayo (2009). However, we abstract away from the choice of identity group – which is often a key focus in this literature. This is both in the interest of tractability and because it is not needed to explain the stylised facts we focus on.
@@ 2 Model
[…]
@@ 2.2 Discussion
Note that in line with typical social identity models, agents dislike being away from the reference point here – either above or below. This takes the view that people want to conform. It is motivated by significant evidence from social psychology, especially regarding descriptive norms (Cialdini et al., 1990; Cialdini and Goldstein, 2004).
This contrasts with the standard assumption in the social comparisons literature that people want to get ahead – and so dislike being below the reference point, but enjoy being above it. Nevertheless, it is possible to replace the social identity style desire to conform with a social comparisons style desire to get ahead. This change does not alter the key insights or the qualitative features of equilibrium behaviour. We show this in the Online Appendix. [^11] To remain highly tractable, it requires a linear, rather than quadratic, cost of consumption.
@@ 4 The effects of changes in income
    In the US [resp. India], people consume less visible goods when the average income of others of the same race in their state [others of the same caste/religious group in their district] is higher (Charles et al., 2009; Khamis et al., 2012).
@@ 5.3 Explaining an empirical feature
We can rationalise these qualitatively different responses to changing inequality across racial groups as perhaps being due to a non-monotonic relationship between community density and community income. More precisely, suppose that density (a la 4) follows a U-shaped pattern: it is first decreasing and then increasing in income. Note that density relates to how people set the ‘benchmark for normal’ that they want to conform to, and hence how strongly people make social comparisons. This non-monotonicity essentially assumes that low and high income agents/communities feel the strongest pressures to conform. It is important to note that this relationship is not causal. Density is determined directly by network structure, not by income. Our suggested relationship may nevertheless be driven by some third factor.
Under this assumption, an increase in inequality due to a transfer that affects only high income communities will involve shifting income from lower density communities to higher density ones. As the resulting changes in incomes are due to a transfer, there will be no knock-on changes to density. 3(iii) tells us that consumption in the communities whose income did not change will then rise. Conversely, an increase in inequality among only low income communities will involve shifting income from higher density communities to lower density ones. In this case, 3(iii) tells us that consumption in the communities whose income did not change will then fall. Figure 1 provides an illustration.
@@
If such a non-monotonic relationship between density and income exists, then we should reasonably expect Blacks to be disproportionately below the trough of the curve, and Whites to be disproportionately above it. This is simply due to the well-documented fact that Blacks have, on average, lower incomes than White (Charles et al., 2009). So an increase in inequality among White would raise consumption of visible goods by Whites (among those whose income did not change), and the reverse would happen for Blacks. This matches exactly the finding by Charles et al. (2009).
Generating the non-monotonic relationship.
We now show how this non-monotonic relationship could be generated by a pair of known facts about the structure of social networks. First, Bailey et al. (2020) show that higher income agents have more geographically dispersed social networks – proportionately less of their friends are nearby. Second, Chetty et al. (2022) show that there is substantial homophily based on socioeconomic status (which is closely tied to income) – people are disproportionately friends with those who have similar income to themselves. Importantly, they find that this homophily is substantially stronger at the upper end of the income distribution.